Interest Rates as Required Rates of Re ...
The time value of money is a concept that states that cash received... Read More
A cumulative distribution offers a convenient tool for determining probabilities for a given random variable. As you have already learnt in a previous learning outcome statement, a cumulative distribution function, F(x), gives the probability that the random variable X is less than or equal to x for every value x. It is usually expressed as:
$$ F(x) = P(X \le x) $$
The random variable X has the following probability distribution function:
$$ \begin{matrix} P(x) = \frac { x }{ 150 } & \text{ for x} = 10, 20, 30, 40, 50 \\ 0 & \text{otherwise} \end{matrix} $$
Calculate and interpret F(20) and F(40), giving an interpretation for each.
Solution
As you will recall, we can determine the probability of each outcome for a random variable given the probability distribution function (pdf).
$$ \begin{align*} P(x) & =\cfrac {x}{150} \\ P(x) & = P(X = x) \\ \end{align*} $$
Therefore,
$$ P(10) =\cfrac {10}{150} $$
Similarly,
$$ P(20) =\cfrac {20}{150} $$
$$ P(30) =\cfrac {30}{150} $$
$$ P(40) =\cfrac {40}{150} $$
And lastly,
$$ P(50) =\cfrac {50}{150} $$
Note: we can prove that our pdf is correct by testing the first rule of probability distribution functions. To do this, we add all the probabilities.
Now,
$$ F(x) = P(X \le x) $$
Therefore,
$$ \begin{align*} F(2) & = P(X \le 20) \\ & = P(X = 10) + P(X = 30) \\ & =\cfrac {10}{150} + \cfrac {20}{150} \\ &=\cfrac {30}{150} \text { or } \cfrac {1}{5} \\ \end{align*} $$
Interpretation: There is a 20% cumulative probability that outcomes 10 or 20 occur.
Similarly,
$$ \begin{align*} F(40) & = P(X \le 40) \\ & = P(X = 10) + P(X = 20) + P(X = 30) + P(X = 40) \\ & =\cfrac {10}{150} + \cfrac {20}{150} + \cfrac {30}{150} + \cfrac {40}{150} \\ & = \cfrac {100}{150} \text{ or } 66.67\% \\ \end{align*} $$
Interpretation: There is a 66.67% cumulative probability that outcomes 10, 20, 30, or 40 occur.
Variable X can take the values 1, 2, 3, and 4. The cumulative probability distribution has been given below. Use it to calculate:
(a) P(X = 2)
(b) P(X = 4)
$$ \begin{array}{c|c|c|c|c} \text{Outcome} & {1} & {2} & {3} & {4} \\ \hline \text{Cumulative Probability distribution} & {0.2} & {0.5} & {0.85} & {1} \\ \end{array} $$
Solution
$$ F(x) = P(X \le x) $$
(a)
$$ \begin{align*} F(2) & = P(X \le 2) = 0.5 \\ 0.5 & = P(X = 1) + P(X = 2) \\ & = 0.2 + P(X = 2)\\ P(X = 2)& = 0.5 – 0.2 = 0.3 \\ \end{align*} $$
Note: a simpler, more direct approach can be:
$$ P(X = 2) = F(2) – F(3) $$
Therefore,
$$ P(X = 2) = 0.5 – 0.2 = 0.3 $$
(b)
$$ \begin{align*} P(X = 4) & = F(4) – F(3) \\ &= 1 – 0.85 = 0.15 \\ \end{align*} $$
Question
Given the following cumulative probability distribution, determine P(2).
$$ \begin{array}{c|c|c|c|c} \text{Outcome} & {0} & {1} & {2} & {3} \\ \hline \text{Cumulative prob.} & {1/8} & {4/8} & {7/8} & {1} \\ \end{array} $$
A. 7/8
B. 3/8
C. 1/8
Solution
The correct answer is B.
$$ \begin{align*} P(2) & = P(X = 2) \\ & = F(2) – F(1) \\ &=\cfrac {7}{8} – \cfrac {4}{8} \\ & =\cfrac {3}{8} \\ \end{align*} $$
Reading 9 LOS 9d:
Calculate and interpret probabilities for a random variable, given its cumulative distribution function.
Master CFA Level I Quantitative Methods concepts, including cumulative distribution functions, probability calculations, random variables, and statistical analysis with study notes, practice questions, and video lessons.
Get Ahead on Your Study Prep This Cyber Monday! Save 35% on all CFA® and FRM® Unlimited Packages. Use code CYBERMONDAY at checkout. Offer ends Dec 1st.